# Package Bids in Combinatorial Electricity Auctions: Selection, Welfare Losses, and Alternatives

### v2

@@ Abstract
A key challenge in combinatorial auctions is designing bid formats that accurately capture agents’ preferences while remaining computationally feasible. This is especially true for electricity auctions, where complex preferences complicate straightforward solutions. In this context, we examine the XOR package bid, the default choice in combinatorial auctions and adopted in European day-ahead and intraday auctions under the name “exclusive group of block bids”. Unlike parametric bid formats often employed in US power auctions, XOR package bids are technology-agnostic, making them particularly suitable for emerging demand-side participants. However, the challenge with package bids is that auctioneers must limit their number to maintain computational feasibility. As a result, agents are constrained in expressing their preferences, potentially lowering their surplus and reducing overall welfare. To address this issue, we propose decision support algorithms that optimize package bid selection, evaluate welfare losses resulting from bid limitations, and explore alternative bid formats. Our findings offer actionable insights for both auctioneers and bidders.
@@ 1 Introduction
However, the two ideal conditions - (i) a perfectly competitive market and (ii) the existence of a Walrasian equilibrium - rarely hold in practice. In the first case, this discrepancy requires the use of market power mitigation methods (Adelowo and Bohland 2024). For the second case, the Walrasian mechanisms employed can only approximate a Walrasian equilibrium, where not every agent is allocated their most profitable bid. While the European algorithm EUPHEMIA does not provide side payments to agents with such “paradoxically” outcomes (NEMO Committee 2024b), various forms of side payments are commonly incorporated into the mechanisms used by auctioneers in the United States (Stevens et al. 2024).
In addressing the above three questions on XOR bids, we assume those two ideals are true: (i) Walrasian equilibria exist, and (ii) no agent has market power. Although we acknowledge the severity of these assumptions, discussions with industry representatives involved in the European auction suggest that they are not uncommon. These assumptions are also made in Karasavvidis et al. (2024), seemingly the only study that addresses XOR bidding in electricity auctions.77endnote: 7Studies on XOR bid selection in other domains include Scheffel et al. (2012), which examines bounded rationality in spectrum auctions, or Hammami et al. (2021), which addresses the “bid construction problem” in transportation auctions. However, we did not find any direct transferable approach to electricity auctions. A key difference might be the high-quality price information available in daily electricity auctions. Furthermore, distinct auction mechanisms, such as pay-as-bid or multi-round formats, require specialized decision support strategies (Pekeč and Rothkopf 2003, Adomavicius et al. 2022). Combinatorial auctions are typically domain specific and often require customized solutions (Bichler et al. 2023). In their work, the authors formulate the XOR bid selection problem as a stochastic bilevel optimization model, where uniform prices are given as exogenous scenarios, and the lower level optimization reflects the bid acceptance rule under the assumed Walrasian equilibrium.88endnote: 8Karasavvidis et al. (2024) do not mention the possibility that this rule may not always hold, but instead implicitly assume equilibrium existence. Industry representatives note that paradoxically rejected or accepted bids are rarely an issue in the European day-ahead market. Although strong evidence for the frequent occurrence of Walrasian equilibria is lacking, some studies, such as a footnote in Reguant (2014) and a simulation by Meeus et al. (2009), suggest that this may be the case.
@@
It is crucial to highlight that day-ahead and intraday auctions are forward markets, requiring participants to account for opportunity costs and risk preferences in their bidding strategies. Opportunity costs may arise naturally within the market or from design flaws, as exemplified by INC-DEC gaming strategies (Graf et al. 2023). In this work, we do not address the estimation of opportunity costs or the handling of risk preferences. Instead, we assume risk-neutral participants who aim to maximize the expected profit based on a generic cost or valuation function1212endnote: 12A body of literature focuses on estimating opportunity costs, particularly in the context of hydropower; for further references, see Löhndorf and Wozabal (2023). Lastly, while extending our methods to convex risk measures, such as conditional value at risk, is conceptually straightforward, this is beyond the scope of the present work.
@@
In general, equilibrium prices cannot be guaranteed in nonconvex markets, though near-equilibrium prices can often be approximated (see, e.g., Starr (1969) or Bichler et al. (2018)). However, unlike electricity auctions, agents in most other auctions typically lack prior knowledge of these prices. Consequently, studies addressing the missing bids problem - such as Goetzendorff et al. (2015) for TV ads and Bichler et al. (2023) for spectrum auctions - advocate for parametric bids, which work as follows: The agent submits parameters $\theta_{1},\ldots,\theta_{n}$, which the auctioneer uses to derive a set $\mathcal{M}(\theta_{1},\ldots,\theta_{n})$ of packages the agent is interested in, along with a function $f(\theta_{1},\ldots,\theta_{n}):\mathcal{M}(\theta_{1},\ldots,\theta_{n})\to\mathbb{R}$ that assigns bid prices to these packages, eliminating the need to enumerate all possible packages. These bids, commonly used for thermal plants and battery storage in US power systems, are known as “multi-part bids” and typically include parameters such as start-up costs or storage capacity (Herrero et al. 2020). A more direct approach would allow agents to submit sets $\mathcal{M}$ and functions $f$ directly, bypassing intermediate parameters. However, the auctioneer must restrict the complexity of $f$ and $\mathcal{M}$ to ensure tractable optimization, making this approach suitable only for agents with sufficiently simple preferences.1414endnote: 14The proposal by Bobo et al. (2021) suggests allowing agents to bid “state variables” and matrices defining a polyhedral set $\mathcal{M}$ and a convex piecewise-linear function $f$. This effectively allows any LP-representable $\mathcal{M}$ and $f$. To handle nonconvexities, the framework could extend to MILP-representable $\mathcal{M}$ and $f$.
@@ 2.5 Infinitely Many Packages
We have identified two cases where (2.1) can be easily solved: when the number of bids allowed exceeds the number of scenarios, or when the agent is only interested in a finite set of packages. However, these conditions may not always hold, as electricity is a divisible good and thus agents could have an infinite number of packages they may be interested in.
@@ 4.1 Case Studies
The plant can generate power within its minimum stable and maximum generation capacities and sell the output in the day-ahead auction. It can ramp up and down between hours, subject to a maximum ramping rate. Production costs consist of marginal costs, start-up and shutdown costs, and no-load costs. Additionally, when the plant is shut down (or started up), it must remain off (or on) for a predefined number of hours before it can start up (or shut down) again. A detailed model is provided in Appendix Appendix B, and the associated parameters are listed in Table 1(a).
@@
The battery can charge and discharge power, allowing it to both buy and sell in the day-ahead auction. Its state of charge must stay within its limits, and both charging/discharging efficiencies are considered. We assume that the battery’s state of charge at the end of the day must be the same as at the beginning of the day. The detailed model is given in Appendix Appendix C, with the parameters listed in Table 1(c).
@@
We consider a utility that needs to supply a heat load using electric boilers, thus purchasing power in the day-ahead auction. The utility has aggregated heat storage available to shift the load between hours. It can also use a gas boiler if electricity prices are too high or reduce the load based on customer contracts, though it is not compensated for any curtailed load. The detailed model is provided in Appendix Appendix D, with the parameters listed in Table 1(b).
@@ 6 Conclusions
This study assumes a competitive equilibrium and does not address price manipulation through market power (Reguant 2014) or exploitation of imperfect equilibrium outcomes (Byers and Eldridge 2023). Theoretical results by Starr (1969) and experiments such as those of Bichler et al. (2018) suggest that while uniform prices may not perfectly align with competitive equilibrium outcomes, they are generally also not too far off. Further research could help to better understand the welfare losses associated with imperfect bid formats when competitive equilibrium is not assumed.
@@ Acknowledgements
Appendix Appendix B Thermal Generator
@@
Appendix Appendix C Battery Storage System
@@
Appendix Appendix D District Heating Utility

### v3

@@ Abstract
A key challenge in combinatorial auctions is designing bid formats that accurately capture agents’ preferences while remaining computationally feasible. This is especially true for electricity auctions, where complex preferences complicate straightforward solutions. In this context, we examine the XOR package bid, the default choice in combinatorial auctions and adopted in European day-ahead and intraday auctions under the name “exclusive group of block bids”. Unlike parametric bid formats often employed in US power auctions, XOR package bids are technology-agnostic, making them particularly suitable for emerging demand-side participants. However, the challenge with package bids is that auctioneers must limit their number to maintain computational feasibility. As a result, agents are constrained in expressing their preferences, potentially lowering their surplus and reducing overall welfare. To address this issue, we propose decision support algorithms that optimize package bid selection, evaluate welfare losses resulting from bid limits, and explore alternative bid formats. In our analysis, we leverage the fact that electricity prices are often fairly predictable and, at least in European auctions, tend to approximate equilibrium prices reasonably well. Our findings offer actionable insights for both auctioneers and bidders.
@@ 1 Introduction
However, the two ideal conditions - (i) a perfectly competitive market and (ii) the existence of a Walrasian equilibrium - cannot be guaranteed. In the first case, this discrepancy requires the use of market power mitigation methods (Adelowo and Bohland 2024). For the second case, the Walrasian mechanisms employed can only approximate a Walrasian equilibrium, where not every agent is allocated their most profitable bid. While the European algorithm EUPHEMIA does not provide side payments to agents with such “paradoxical” outcomes (NEMO Committee 2024b), various forms of side payments are commonly incorporated into the mechanisms used by auctioneers in the United States (Stevens et al. 2024).
In addressing the above three questions on XOR bids, we assume those two ideals are true: (i) Walrasian equilibria exist, and (ii) no agent has market power. While these assumptions are admittedly strong, discussions with industry stakeholders involved in European auctions - along with supporting evidence from Hübner (2025) and Graf and Wozabal (2013) - suggest that such assumptions are common in practice and approximate real-world conditions reasonably well. Notably, the study by Karasavvidis et al. (2024) - to the best of our knowledge the only one to address XOR bidding in electricity auctions - also adopts these assumptions.77endnote: 7Studies on XOR bid selection in other domains include Scheffel et al. (2012), which examines bounded rationality in spectrum auctions, or Hammami et al. (2021), which addresses the “bid construction problem” in transportation auctions. However, we did not find any direct transferable approach to electricity auctions. A key difference might be the high-quality price information available in daily electricity auctions. Furthermore, distinct auction mechanisms, such as pay-as-bid or multi-round formats, require specialized decision support strategies (Pekeč and Rothkopf 2003, Adomavicius et al. 2022). Combinatorial auctions are typically domain specific and often require customized solutions (Bichler et al. 2023). Most studies on bidding in electricity auctions focus on hourly bids (see the review by Aasgård et al. 2019), while block bids have received comparatively less attention (e.g., Fleten and Kristoffersen 2007, Karasavvidis et al. 2021). To the best of our knowledge, Karasavvidis et al. (2024) is the only study that explicitly examines XOR bids in electricity auctions. In their work, Karasavvidis et al. (2024) formulate the XOR bid selection problem as a stochastic bilevel optimization model, where uniform prices are given as exogenous scenarios, and the lower level optimization reflects the bid acceptance rule under the assumed Walrasian equilibrium.88endnote: 8Karasavvidis et al. (2024) do not mention the possibility that this rule may not always hold, but instead implicitly assume equilibrium existence. Industry representatives note that paradoxically rejected or accepted bids are rarely an issue in the European day-ahead market. Indeed, in 2023, an average of 4,401 block bids were submitted per day, while only 7.5 were paradoxically rejected (NEMO Committee 2024a). A more comprehensive study on this phenomenon can be found in Hübner (2025).
@@
It is crucial to highlight that day-ahead and intraday auctions are forward markets, requiring participants to account for opportunity costs and risk preferences in their bidding strategies. Opportunity costs may arise naturally within the market or from design flaws, as exemplified by INC-DEC gaming strategies (Graf et al. 2023). In this work, we do not address the estimation of opportunity costs or the handling of risk preferences. Instead, we assume risk-neutral participants who aim to maximize the expected profit based on a generic cost or valuation function1212endnote: 12A body of literature focuses on estimating opportunity costs, particularly in the context of hydropower; for further references, see Löhndorf and Wozabal (2023). In Appendix Appendix C, we study risk-averse participants maximizing the conditional value at risk (CVaR) .
@@
In general, equilibrium prices cannot be guaranteed in nonconvex markets, though near-equilibrium prices can often be approximated (see, e.g., Starr (1969), Bichler et al. (2018), or Hübner (2025)). However, unlike electricity auctions, agents in most other auctions typically lack prior knowledge of these prices. Consequently, studies addressing the missing bids problem - such as Goetzendorff et al. (2015) for TV ads and Bichler et al. (2023) for spectrum auctions - advocate for parametric bids, which work as follows: The agent submits parameters $\theta_{1},\ldots,\theta_{n}$, which the auctioneer uses to derive a set $\mathcal{M}(\theta_{1},\ldots,\theta_{n})$ of packages the agent is interested in, along with a function $f(\theta_{1},\ldots,\theta_{n}):\mathcal{M}(\theta_{1},\ldots,\theta_{n})\to\mathbb{R}$ that assigns bid prices to these packages, eliminating the need to enumerate all possible packages. These bids, commonly used for thermal plants and battery storage in US power systems, are known as “multi-part bids” and typically include parameters such as start-up costs or storage capacity (Herrero et al. 2020). A more direct approach would allow agents to submit sets $\mathcal{M}$ and functions $f$ directly, bypassing intermediate parameters. However, the auctioneer must restrict the complexity of $f$ and $\mathcal{M}$ to ensure tractable optimization, making this approach suitable only for agents with sufficiently simple preferences.1414endnote: 14The proposal by Bobo et al. (2021) suggests allowing agents to bid “state variables” and matrices defining a polyhedral set $\mathcal{M}$ and a convex piecewise-linear function $f$. This effectively allows any LP-representable $\mathcal{M}$ and $f$. To handle nonconvexities, the framework could extend to MILP-representable $\mathcal{M}$ and $f$.
@@ 2.4 Finitely Many Packages
In Appendix Appendix B, we highlight that our package bid selection problem (2.4) is a special variant of the bilevel knapsack problem. Interestingly, Proposition 2 shows that this specific variant can be solved in polynomial time. This is surprising, since - as discussed by Caprara et al. (2014) - classical knapsack problems are NP-complete, and bilevel problems built on top of NP-complete single-level problems are typically $\Sigma_{2}^{p}$-complete. This implies that they cannot be reformulated as single-level integer programs of polynomial size unless the polynomial hierarchy collapses. Moreover, Caprara et al. (2014) showed that many bilevel knapsack variants studied in the literature are indeed $\Sigma_{2}^{p}$-complete. Our variant is, therefore, exceptionally tractable.
While we assumed a risk-neutral agent maximizing expected surplus, Proposition 2 remains valid as long as the agent optimizes an LP-representable risk measure like, for example, the Conditional Value at Risk (CVaR) (Rockafellar and Uryasev 2000). Appendix Appendix C provides a CVaR-based formulation, which, in our case, maximizes the expected worst-case profit.
@@ 2.5 Infinitely Many Packages
We have identified two cases where (2.1) can be easily solved: when the number of bids allowed exceeds the number of scenarios, or when the agent is only interested in a finite set of packages. However, these conditions may not always hold, as electricity is a divisible good, and thus, agents could have an infinite number of packages they may be interested in.
@@ 4.1 Case Studies
The plant can generate power within its minimum stable and maximum generation capacities and sell the output in the day-ahead auction. It can ramp up and down between hours, subject to a maximum ramping rate. Production costs consist of marginal costs, start-up and shutdown costs, and no-load costs. Additionally, when the plant is shut down (or started up), it must remain off (or on) for a predefined number of hours before it can start up (or shut down) again. A detailed model is provided in Appendix Appendix D, and the associated parameters are listed in Table 1(a).
@@
The battery can charge and discharge power, allowing it to both buy and sell in the day-ahead auction. Its state of charge must stay within its limits, and both charging/discharging efficiencies are considered. We assume that the battery’s state of charge at the end of the day must be the same as at the beginning of the day. The detailed model is given in Appendix Appendix E, with the parameters listed in Table 1(c).
@@
We consider a utility that needs to supply a heat load using electric boilers, thus purchasing power in the day-ahead auction. The utility has aggregated heat storage available to shift the load between hours. It can also use a gas boiler if electricity prices are too high or reduce the load based on customer contracts, though it is not compensated for any curtailed load. The detailed model is provided in Appendix Appendix F, with the parameters listed in Table 1(b).
@@ 6 Conclusions
This study assumes a competitive equilibrium and does not address price manipulation through market power (Reguant 2014) or exploitation of imperfect equilibrium outcomes (Byers and Eldridge 2023). Theoretical and empirical evidence by Starr (1969)Bichler et al. (2018) or Hübner (2025) suggest that while uniform prices may not perfectly align with competitive equilibrium outcomes, they are often also not too far off. Further research could help to better understand the welfare losses associated with imperfect bid formats when competitive equilibrium is not assumed.
@@ Acknowledgements
Appendix Appendix B Package Bid Selection as Bilevel Knapsack Problem
The package bid selection problem (2.4), along with its reformulations (12) and (2.4), can be equivalently expressed as the following bilevel knapsack problem:
$\max_{\delta}$ $\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\sum_{k\in\mathcal{K}}U_{ks}\cdot\gamma_{ks}$ (26a)
s.t. $\quad\sum_{k\in\mathcal{K}}\delta_{k}\leq B$ (26b)
$\quad\delta_{k}\in\{0,1\}\quad\forall k\in\mathcal{K}$ (26c)
$\quad\big{(}\gamma_{ks}\big{)}_{k\in\mathcal{K},s\in\mathcal{S}}\in$ (26d)
$\quad\operatorname*{arg\,max}_{\gamma}\quad\sum_{s\in\mathcal{S}}\pi_{s}\cdot\sum_{k\in\mathcal{K}}U_{ks}\cdot\gamma_{ks}$ (26e)
[…]
$\qquad\qquad\quad\sum_{k\in\mathcal{K}}\gamma_{ks}\leq 1\quad\forall s\in\mathcal{S}$ (26g)
$\qquad\qquad\quad\gamma_{ks}\in\{0,1\}\quad\forall s\in\mathcal{S},\;k\in\mathcal{K}.$ (26h)
where $U_{ks}=\bar{v}_{k}-\langle\lambda_{s},\bar{x}_{k}\rangle$ represents the utility of selecting package $k$ in scenario $s$. The leader (26a)-(26d) (“bidder”) selects up to $B$ packages to include in the knapsack. The follower (26e)-(26h) (“auctioneer”) can then take one package out of the knapsack for each scenario $s\in\mathcal{S}$. The leader and follower have the same objective; they both try to maximize the expected profit of the selected packages.
This bilevel knapsack problem is especially tractable since the leader and follower share the same objective, causing the bilevel structure to collapse into a single-level problem with a totally unimodular constraint matrix (Proposition 2).
Appendix Appendix C Optimizing the Conditional Value at Risk
In formulating the bid selection problem (2.1), we assumed a risk-neutral agent maximizing expected profit. However, risk-averse agents may prefer to optimize expected profit in the worst $1-\beta\%$ of cases - that is, to maximize the $\beta$-CVaR. Using the well-known reformulation by Rockafellar and Uryasev (2000), this can be expressed as:
$\max_{x_{b},p_{b},\alpha,\eta_{s}}$ $\quad\alpha-\frac{1}{1-\beta}\cdot\sum_{s\in\mathcal{S}}\pi_{s}\cdot\eta_{s}$
s.t. $\quad\eta_{s}\geq\alpha-v(x_{s}^{\ast})+\langle\lambda_{s},x_{s}^{\ast}\rangle\quad\forall s\in\mathcal{S}$
$\quad\eta_{s}\geq 0,\;\alpha\in\mathbb{R}\quad\forall s\in\mathcal{S}$
[…]
Theorem 1 in Rockafellar and Uryasev (2000) shows that the objective function $\alpha-\frac{1}{1-\beta}\sum_{s\in\mathcal{S}}\pi_{s}\cdot\eta_{s}$ corresponds to the $\beta$-CVaR, with $\alpha$ representing the $\beta$-VaR at the optimum. Theorem 2 confirms the correctness of the reformulation.
Appendix Appendix D Thermal Generator
@@
Appendix Appendix E Battery Storage System
@@
Appendix Appendix F District Heating Utility